# Non-Bloch band theory for time-modulated discrete mechanical systems

### v1

@@ 3.2 Non-Bloch band theory with generalized Brillouin zone
We numerically show that this non-Bloch band theory predicts the spectrum of the finite-length time-modulated system. We use the same parameters and compute $\beta^{(1)}=\beta_{N}$ and $\beta^{(2)}=\beta_{N+1}$ for varying $\omega\in\mathbb{C}$. The results are plotted in Figure 5 (a) and indicate that the contour $|\beta^{(1)}(\omega)|=|\beta^{(2)}(\omega)|$ (i.e. $\sigma_{\mathrm{GBZ}}$) forms a continuous curve with several branches in the complex $\omega$-plane, which resembles the finite-chain spectrum $\sigma_{\mathrm{OBC}}$ in Figure 3 (c). In fact, Figure 5 (b) shows that almost all the eigenvalues in $\sigma_{\mathrm{OBC}}$ lie on the contour $\sigma_{\mathrm{GBZ}}$. Figure 5 (d) shows $C_{\mathrm{GBZ}}$ calculated from the spectrum $\sigma_{\mathrm{GBZ}}$. In contrast to the standard BZ (unit circle $|\beta|=1$), the GBZ deviates from the circle, which suggests the existence of non-Hermitian skin effects [25]. An eigenmode corresponding to an eigenvalue $\omega\in\sigma_{\mathrm{GBZ}}$ is called a non-Hermitian skin mode if $|\beta_{N}(\omega)|=|\beta_{N+1}(\omega)|\neq 1$. For example, the eigenvalue A in Figure 5 (b) lies on the spectrum $\sigma_{\mathrm{GBZ}}$ with $|\beta|>1$. This implies that the corresponding eigenmode grows exponentially as the site index $j$ increases. In fact, the finite-length chain exhibits a growing eigenmode at A as we have seen in Figure 3 (d). The same discussion is valid for other eigenvalues B, C, D, E, and their conjugate pairs, i.e., the non-Bloch band theory with the GBZ tells whether a given eigenvalue is associated with a mode localized at the edges (non-Hermitian skin mode) or extending in the bulk (usual Bloch mode).
@@ 3.3 Transient wave motion
with the angular frequency $\omega_{\mathrm{f}}=0.31\omega_{0}$ and constant amplitude $A>0$.
@@
$N-1$ $5.564\times 10^{-1}-6.880\times 10^{-1}\mathrm{i}$ $-8.908\times 10^{-1}+1.223\times 10^{-1}\mathrm{i}$
$N$ $5.748\times 10^{-1}+8.183\times 10^{-1}\mathrm{i}$ $9.584\times 10^{-1}+2.656\times 10^{-14}\mathrm{i}$
@@
$N+2$ $7.106\times 10^{-1}-8.787\times 10^{-1}\mathrm{i}$ $-8.908\times 10^{-1}-1.223\times 10^{-1}\mathrm{i}$

### v2

@@ 3.2 Non-Bloch band theory with generalized Brillouin zone
We numerically show that this non-Bloch band theory predicts the spectrum of the finite-length time-modulated system. We use the same parameters and compute $\beta^{(1)}=\beta_{N}$ and $\beta^{(2)}=\beta_{N+1}$ for varying $\omega\in\mathbb{C}$. The results are plotted in Figure 5 (a) and indicate that the contour $|\beta^{(1)}(\omega)|=|\beta^{(2)}(\omega)|$ (i.e. $\sigma_{\mathrm{GBZ}}$) forms a continuous curve with several branches in the complex $\omega$-plane, which resembles the finite-chain spectrum $\sigma_{\mathrm{OBC}}$ in Figure 3 (c). In fact, Figure 5 (b) shows that almost all the eigenvalues in $\sigma_{\mathrm{OBC}}$ lie on the contour $\sigma_{\mathrm{GBZ}}$. Figure 5 (d) shows $C_{\mathrm{GBZ}}$ calculated from the spectrum $\sigma_{\mathrm{GBZ}}$. In contrast to the standard BZ (unit circle $|\beta|=1$), the GBZ deviates from the circle, which suggests the existence of non-Hermitian skin effects yao2018edge . An eigenmode corresponding to an eigenvalue $\omega\in\sigma_{\mathrm{GBZ}}$ is called a non-Hermitian skin mode if $|\beta_{N}(\omega)|=|\beta_{N+1}(\omega)|\neq 1$. For example, the eigenvalue A in Figure 5 (b) lies on the spectrum $\sigma_{\mathrm{GBZ}}$ with $|\beta|>1$. This implies that the corresponding eigenmode grows exponentially as the site index $j$ increases. In fact, the finite-length chain exhibits a growing eigenmode at A as we have seen in Figure 3 (d). The same discussion is valid for other eigenvalues B, C, D, E, and their conjugate pairs, i.e., the non-Bloch band theory with the GBZ tells whether a given eigenvalue is associated with a mode localized at the edges (non-Hermitian skin mode) or extending in the bulk (usual Bloch mode).
@@
The existence of non-Hermitian skin modes is closely related to the reciprocity of a system. The periodic chain with three mass-spring pairs is spatially asymmetric due to the phase differences of stiffness modulation. This asymmetry is necessary for breaking reciprocity. Reciprocity implies that eigenfrequencies $\omega$ are invariant with respect to the wavenumber inversion $k\to-k$, meaning that each eigenfrequency trajectory does not form a closed loop encircling a nonzero area when $k$ varies along the BZ (from $\pi$ to $\pi$). This concludes that the non-Hermitian skin effect does not manifest itself lin2023topological . In fact, when the modulation phase $\phi_{3}$ is changed as $\phi_{3}=0$, i.e., $\phi_{1}=\phi_{3}$, the system is no longer non-reciprocal ammari2022nonreciprocal . In this case, we numerically observed that the GBZ coincides with the standard BZ (unit circle $|\beta|=1$ in the complex $\beta$-plane) while the corresponding quasifrequency spectrum comprises complex numbers with positive and negative imaginary parts. From this observation, we conclude that both the spatial asymmetry and stiffness modulation are essential for achieving the non-Hermitian skin effect, characterized by the discrepancy between the standard BZ and GBZ. In such cases, the proposed GBZ theory gives useful and quantitative information about non-Hermitian skin modes.
@@ 3.3 Transient wave motion
with the angular frequency $\omega_{\mathrm{f}}=0.31\omega_{0}$ and constant amplitude $A>0$. Note that the phase of the forcing does not affect the amplitudes of excited modes.
@@
N1N-1 $5.564\times 10^{-1}-6.880\times 10^{-1}\mathrm{i}$ $8.908\times 10^{-1}+1.223\times 10^{-1}\mathrm{i}$
NN $5.748\times 10^{-1}+8.183\times 10^{-1}\mathrm{i}$ $9.584\times 10^{-1}+2.656\times 10^{-14}\mathrm{i}$
@@
$N+2$ $7.106\times 10^{-1}-8.787\times 10^{-1}\mathrm{i}$ $8.908\times 10^{-1}-1.223\times 10^{-1}\mathrm{i}$
@@ 4 Concluding remarks
As we focused on eigenvalues and eigenmodes of Floquet systems without forcing, a future direction of this work is a more detailed analysis of the excitation of Bloch or non-Hermitian skin modes in response to given forces. For example, we are interested in the relationship between excited amplitudes and modulation or forcing phases, which is well studied in single resonator models described by the Mathieu equation acar2016floquet . In addition, higher-dimensional lattices are of interest as such systems can possibly exhibit edge/corner skin effects or unidirectional mode conversion.
We note that other mathematical tools may also be useful for the analysis of time-modulated systems. For instance, a higher-order homogenization method for the wave equation with time-varying coefficients was recently proposed in touboul2024high . It is worth discussing whether such homogenization-based approaches are valid for time-modulated and non-reciprocal models exhibiting the non-Hermitian skin effect.